freezer has a coefficient of performance of 2.40. The freezer is to convert 1.80 kg of water at 25.0°C to

1.80 kg of ice at -5.0°C in one hour. (a) What amount of heat must be removed from the water at 25.0°C

to convert it to ice at -5.0°C? (b) How much electrical energy is consumed by the freezer during this

hour? (c) How much wasted heat is delivered to the room in which the freezer sits?


1
Expert's answer
2021-05-16T17:58:13-0400

"COP = 2.40
\\ m_i = 1.80kg\\
T_1= 25.0°C
\\ T_2 = -5.0°C\\
t = 1\ hr = 3600\ sec"


"(a)\ Q = mc∆T_1+ mL + mc∆T_2\\
Q = m(c∆T_1+ L + c∆T_2)\\\\
Q = 1.80(4190(25) + 334000 + 4190(5))= 808000\ J"


"808\ kJ" of heat must be removed from the water.



"(b)\ COP = \dfrac{Q_c}W"

"2.40 = \dfrac{808}{W}"


"W= 337\ kJ"



"(c)\ Q_w= W+ Q_c\\

Q_w = 337 + 808=1145k J"


"\therefore 1145\ kJ" wasted heat is delivered to the room in which the freezer sits